OpenStudy (anonymous):

Simplify (5x+3) - 2(2x-4).

OpenStudy (anonymous):

distribute the -2

OpenStudy (anonymous):

so (5x+3 first so all laid out its 5*x+3 so do 5-3 you get what x is so tell me what x is

OpenStudy (princeharryyy):

x + 11

OpenStudy (princeharryyy):

@jenise

OpenStudy (anonymous):

so (5x+3 first so all laid out its 5*x+3 so do 5-3 you get what x. x=2 so 5x2+3=13 so than 13-2=11 than do the say with 2x-4 all laid out is 2*x-4 so 4+2=6 so go back to 11x6= and what ever 11x6 is your answer hope i helped

OpenStudy (princeharryyy):

5x+3-4x+8 => x+11 as the answer @helpmeplzzzzz! n @jenise